# Automated threshold selection and associated inference uncertainty for univariate extremes

### v3

@@ 7 Application to river flow data
[…]
Table 7 shows the selected thresholds of each of the methods for a range of candidate grids [^4] In marked cases, the Northrop method outputted a chosen threshold with some convergence warnings. . The remarkable robustness of the EQD (evaluated with $k=200$ bootstrap samples) across grids stems from the method’s novel incorporation of data uncertainty. The Wadsworth method fails to estimate a threshold unless the grid is made very coarse, and even then exhibits considerable sensitivity (varying between 0% and 90% sample quantiles) over grids of equal size but different endpoints and increments. This is problematic as a coarse grid is likely to remove the most appropriate threshold from consideration. The Northrop method critically depends on the validation threshold, and we find that increasing this level above the $90\%$-quantile leads to failure or convergence warnings. The thresholds selected by this method are quite variable (between 0% and 80% sample quantiles) over the grids.
@@
Figure 2 shows a QQ-plot for the GPD model using the EQD estimate $\hat{u}=67.10$. The tolerance bounds show a reasonable agreement between model and data. For $\hat{u}$, Figure 2 also shows the $T$-year return level estimates, with $1\leq T\leq 1000$. The 95% CIs incorporate parameter uncertainty alone and both parameter and threshold uncertainty using Algorithms 1 and 2 respectively, with an increase in uncertainty from the latter for larger $T$; e.g., for the 100- and 1000-year return levels, the CI width increases by a factor of 1.38 and 1.52 respectively. This reiterates how vital it is to incorporate threshold uncertainty into inference.
@@ 8 Conclusion and Discussion
We proposed two substantial developments to univariate extreme value analysis. Firstly, we addressed the widely-studied problem of how to automatically select/estimate a threshold above which an extreme value, generalised Pareto, model can be fitted. We presented a novel and simple approach, which we termed the EQD method. Secondly, we proposed a new approach to improve the calibration of confidence intervals for high quantile inference, addressing an important but under-studied problem. We achieve this through an intuitively simple, but computationally intensive, double-bootstrapping technique which propagates the uncertainty in the threshold estimation through to quantile inference.
Regarding the threshold selection component of the work, we compared the EQD method to the leading existing threshold selection methods in terms of both threshold selection and consequent high quantile estimation. This was conducted using data from iid continuous univariate random variables and the superiority of the EQD method was illustrated across a range of examples using various metrics. Relative to existing approaches, we showed that the EQD exhibits greater robustness to changes in the set of candidate thresholds, to tuning parameters, and avoids a reliance on asymptotic theory in existing likelihood methods.
So why does the EQD method perform much better than the existing approaches? Ultimately, it is because it better accounts for the variability in the resulting GPD parameter estimates. Specifically, this variability is assessed across replicated bootstrapped samples of the same size as the observed data and is incorporated into a summary metric for each of the candidate thresholds. In contrast, the existing methods do not attempt to account for this characteristic, and as a result, overly focus on the fit to the actual sample. Consequently, over repeated samples, the existing methods exhibit more bias and much larger variability in their selected thresholds and estimated quantiles.
@@ Acknowledgements
Conor Murphy, Jonathan A. Tawn
Department of Mathematics and Statistics, Lancaster University
and
Zak Varty
Department of Mathematics, Imperial College London
@@
This document provides further information to accompany the main paper. Section S:2 includes derivations of key properties, further description of methods and additional experiments using some of the standard methods for threshold selection, as referenced in Section 2 of the main paper. Section S:3 provides more detailed descriptions of the Wadsworth, (2016) and Northrop et al., (2017) methods that are used in the analyses of the main paper. Section S:4 presents further details of the distributions in Cases 1-4 of Section 6 of the main paper, including quantile derivations. Section S:5 covers the reasoning for the omission of the Varty et al., (2021) approach from the simulation study in the main paper, justifies the choice of the default tuning parameters for the EQD method and discusses the choice of calibration data within the definition of $d_{b}(u)$ in Section 4 of the main text. Section S:6 presents additional simulation experiments and a detailed breakdown of the results outlined in Section 6 of the main paper. Finally, Section S:7 provides a more extensive analysis of the coverage of true quantiles achieved by both Algorithm 1 and 2, as mentioned in Section 5 of the main text.
@@
$F_{X}(x)=\int_{0}^{x}\frac{h(s;\sigma,\xi)\mathbb{P}(B<s)}{q+\bar{H}(1;\sigma,\xi)}\ \mathrm{d}s$ (S.6)
where $\bar{H}(x;\sigma,\xi)$ and $h(x;\sigma,\xi)$ are the survivor and density functions of a GPD with parameters $(\sigma,\xi)$ and a threshold 00, $B\sim\text{Beta}(\alpha,\beta)$, so $0\leq B\leq 1$ and $q=\int_{0}^{1}h(s;0.5,0.1)\mathbb{P}(B<s)\mathrm{d}s$. This unusual distribution function transitions from a non-GPD distribution to an exact GPD for excesses of $1$ (the upper bound of $B$), so that the true threshold for $X$ is $u=1$ and the excess distribution, for $x>1$, has survivor function $\bar{H}(x-1;\sigma_{1}=\sigma+\xi,\xi)$ following from the threshold stability property (S.1). Here $\tau:=\mathbb{P}(X\leq 1)=q/(q+\bar{H}(1;\sigma,\xi))$ so the true quantile $x_{p}$ for $X$ with exceedance probability $p$, where $p\leq 1-\tau$, is:
@@
Although $F_{X}$ appears complex, it is straightforward to simulate samples from $X$. Specifically, $X$ is generated using a rejection method by first simulating a proposal variable $Y\sim\text{GPD}(\sigma,\xi)$ above the threshold of 00. The rejection step involves rejecting $Y$ as a candidate for $X$ only if $Y<B$, where $B$ is a random variate generated from a $\text{Beta}(\alpha,\beta)$ distribution. Since $B\leq 1$, for $Y>1$ it follows that $X=Y$, but only a proportion of the values of $Y<1$ are retained in $X$, with the rate of retention dependent on the parameters of the $\text{Beta}(\alpha,\beta)$ distribution.
@@
In this section, we provide evidence, based on a range of simulation studies, to support several decisions we made in Section 4 of the main paper. Specifically, in Section S:5.2, we provide evidence to demonstrate the advantages of using the EQD over the Varty et al., (2021) method; Section S:5.3 outlines results indicating that our suggested default choices for the tuning parameters $(k,m)$ of the EQD method are widely suitable, and that there is very little sensitivity to these choices; and in Section S:5.4, we show that the choice of bootstrap data as the calibration data in the metric $d_{b}(u)$ works comparably relative to using the observed data.
@@
We assess the performance of the two methods based on both the selection of thresholds (as the truth is known in each case) and on the subsequent estimation of quantiles for a range of exceedance probabilities, namely the $(1-p_{j})$-quantiles where $p_{j}=1/(10^{j}n)$ for $j=0,1,2$ with $n$ denoting the length of the simulated dataset. For each case and each of the measures of fit, all comparisons between the methods are based on estimates obtained using the same set of 500 replicated samples, so any difference in the methods found is simply due to the two method’s performance.
@@
0 0.190 0.190 $\bm{0.691}$ 0.731
1 $\bm{0.323}$ 0.324 $\bm{1.658}$ 1.765
2 0.483 0.483 $\bm{3.363}$ 3.615
Table S.2: RMSE of the estimated quantiles in Cases 1-4 based on fitted GPD above chosen threshold for the EQD and Varty et al., (2021) methods. The smallest value for each quantile are highlighted in bold.
@@
0 0.036 0.036 $\bm{0.454}$ 0.522
1 0.105 0.105 $\bm{2.622}$ 3.074
2 0.233 0.233 $\bm{10.940}$ 13.000
Table S.3: Bias [left] and variance [right] of the estimated quantiles in Cases 1-4 based on fitted GPD above chosen threshold for the EQD and Varty et al., (2021) methods. The smallest variance and absolute bias for each quantile are highlighted in bold.
@@
This section provides the sources of evidence presented in Section 4.2 of the main paper. Specifically, it details an analysis of the sensitivity of the EQD method to different choices of its two tuning parameters $(k,m)$, where $k$ denotes the number of the bootstraps for which $d_{b}(u)$ is evaluated in order to calculate the overall metric $d_{E}(u)$, and $m$ is the number of quantiles used in the evaluation of the metric $d_{b}(u)$ for each bootstrap. In the main paper, the values of $(k,m)=(100,500)$ are proposed as the default values for the simulation studies of the performance of the EQD method. These values are used for the tuning parameters in all simulation studies of the EQD method in the main paper and the supplementary material.
@@
We focus our sensitivity analysis on the 500 replicated samples of Case 1 where $n=1000$, detailed in Section 6 of the main paper. Tables S.4 & S.5 provide the RMSEs of the threshold estimates along with the computation time relative to that of the default value when using the EQD with different values of $k$ and $m$ respectively.
@@
First we consider the choice of $k$, the number of the bootstraps for which $d_{b}(u)$ is evaluated when calculating $d_{E}(u)$. Table S.4 shows that the computation time of the EQD method increases linearly with $k$. As $k$ is simply the number of bootstrap samples in an average, then in principle we want to take $k$ as large as possible to remove Monte Carlo noise in the average approximation of the associated expectation. Thus, in selecting $k$, we require it to be sufficiently large so that any residual Monte Carlo noise is not important (for threshold selection to be stable) whilst recognising the linear increase in computation time from this choice. Thus, for one-off analyses, as computation time is not of particular concern, it is ideal to take $k$ as large as possible. However, for simulation studies, a more careful choice of $k$ is required as accuracy needs to be balanced with computation time. Table S.4 provides evidence on how the Monte Carlo noise is diminishing as $k$ increases, and shows the RMSE value stabilising as $k$ increases. The decreases in RMSE are only very slight, especially when compared to the differences we see between the EQD and other existing approaches in Section 6 of the main paper.
Now we consider the choice of $m$, the number of quantiles used in the evaluation of the metric $d_{b}(u)$ for each bootstrap. Table S.5 shows results for two different strategies for selecting $m$: the first allowing $m$ to be proportional to the data sample size $n$ irrespective of threshold value, i.e. $m=cn$ for $c=0.5,1,2,10$, and the second allowing $m$ to vary according to the number, $n_{u}$, of exceedances of each candidate threshold $u$, i.e. $m=cn_{u}$ for $c=0.5,1,2,10$. Our reason for exploring the second strategy is to ensure that across candidate thresholds we are using the same level of interpolation/extrapolation to non-sample quantiles. For each strategy we examine the effect of different degrees of proportionality $c$. The RMSE of the threshold estimates obtained using the EQD show little sensitivity to the value of $m$ across the two strategies and all levels of proportionality.
Table S.5 also shows that increasing $m$ in either strategy is essentially wasting the increased computation time. When estimating $d_{b}(u)$ for a particular bootstrap, we are aiming to approximate the integrated absolute error between the model and sample quantiles. This sensitivity analysis shows that once we choose a suitably large value for $m$, any changes in this approximation error are quite small in comparison to the $d_{b}(u)$ value itself. Since $m$ controls the quality of this approximation, we consider it best to keep $m$ fixed across thresholds. We suggest $m=500$ is sufficient based on the comparable quality of results for $m=0.5n$ in this case and the associated limited computation time. Given we are confident in the quality of results for this choice, for convenience, we keep this fixed value across sample sizes and threshold levels in our simulation studies.
In conclusion, in order to balance accuracy and computation time for our large scale simulation studies, we find that the EQD method, with value $k=100$ and $m=500$, provides sufficiently accurate results in a timely manner, so we choose to use these value throughout the paper, unless stated otherwise.
$k$ 200 400 1000 100 (default)
@@
Table S.4: RMSEs for threshold estimates and for the relative computation time compared to the default choice of $k=100$ obtained using the EQD method for different values of $k$ for Case 1. Each result uses $m=500$ and 500 replicated samples.
@@
Table S.5: RMSEs for threshold estimates and for the relative computation time compared to the default of $m=500$ of the EQD method for different values of $m$ for Case 1. Each result uses $k=100$ and 500 replicated samples.
@@
We here provide results for an adjusted version the EQD method (as suggested by a referee) where the calibration data used for threshold estimation are the actual observed sample excesses $\bm{x}_{u}$ of candidate threshold $u$, as described below. This adjusted version differs from the EQD method as proposed in Section 4 of the main paper, where the calibration data for a particular bootstrapped sample are calculated based on the bootstrapped excesses $\bm{x}^{b}_{u}$. Thus, the only difference between the two approaches is that in $d_{b}(u)$, given by metric (2), the $Q(p_{j},\bm{x}^{b}_{u})$ term is replaced by $Q(p_{j},\bm{x}_{u})$ in the adjusted version.

### v4

@@ 7 Application to river flow data
The River Nidd dataset consists of 154 storm event peak daily river flow rates that exceeded 65 m3/s in the period 1934-1969, i.e., an average exceedance rate of 4.4 events per year. Each observation can be deemed “extreme” and iid, though not necessarily well-described by a GPD. Davison and Smith, (1990) identify the difficulties these data present for threshold selection and parameter uncertainty, which we reiterated in discussion of Figure 1. Given the small sample size for the River Nidd, any increase in the threshold value is more significant in terms of parameter uncertainty, than for larger datasets studied in Section 6.
Table 7 shows the selected thresholds of each of the methods for a range of candidate grids [^3] In marked cases, the Northrop method outputted a chosen threshold with some convergence warnings. . The remarkable robustness of the EQD (evaluated with $B=200$ bootstrap samples) across grids stems from the method’s novel incorporation of data uncertainty. The Wadsworth method fails to estimate a threshold unless the grid is made very coarse, and even then exhibits considerable sensitivity (varying between 0% and 90% sample quantiles) over grids of equal size but different endpoints and increments. This is problematic as a coarse grid is likely to remove the most appropriate threshold from consideration. The Northrop method critically depends on the validation threshold, and we find that increasing this level above the $90\%$-quantile leads to failure or convergence warnings. The thresholds selected by this method are quite variable (between 0% and 80% sample quantiles) over the grids.
@@
Figure 2 shows a QQ-plot for the GPD model using the EQD estimate $\hat{u}=67.10$. The tolerance bounds show a reasonable agreement between model and data. For $\hat{u}$, Figure 2 also shows the $T$-year return level estimates, with $1\leq T\leq 1000$. The 95% CIs incorporate parameter uncertainty alone and both parameter and threshold uncertainty via Algorithms 1 and 2 respectively, with an increase in uncertainty from the latter for larger $T$; e.g., for the 100- and 1000-year return levels, the CI width increases by a factor of 1.38 and 1.52 respectively. This reiterates how vital it is to incorporate threshold uncertainty into inference.
@@ 8 Conclusion and discussion
We proposed two substantial developments to univariate extreme value analysis. Firstly, we addressed the widely-studied problem of how to automatically select/estimate a threshold above which an extreme value, generalised Pareto, model can be fitted. We presented a novel and simple approach, which we termed the EQD method, that minimises an approximation to the IAE of the model quantiles and quantiles of the data generating process. Secondly, we proposed a new approach to improve the calibration of confidence intervals for high quantile inference, addressing an important but under-studied problem. We achieve this through an intuitively simple, but computationally intensive, double-bootstrapping technique which propagates the uncertainty in the threshold estimation through to quantile inference.
Regarding the threshold selection component of the work, we compared the EQD method to the leading existing threshold selection methods in terms of both threshold selection and consequent high quantile estimation. This was conducted using data from iid continuous univariate random variables and the superiority of the EQD method was illustrated across a range of examples using various metrics. Relative to existing approaches, we showed that the EQD exhibits greater robustness to changes in the set of candidate thresholds, to tuning parameters, and avoids a reliance on asymptotic theory in existing likelihood methods. The EQD method is applicable for all data set sizes and for any set of candidate thresholds.
So why does the EQD method perform much better than the existing approaches? Our analysis has identified two core reasons: the choice of a robust measure of goodness of fit for a given (bootstrapped) sample, which controls bias; and the use of bootstrapped replicates, which leads to reduced variance and also appears to reduce bias. Specifically, in comparison to existing methods, the use of our goodness-of-fit measure, over simply exploiting the GPD threshold stability property, ensures better model fits and hence better threshold selection, and the bootstrapping removes the variation that arises if only the observed sample is used, as that may not be a typical realisation from the underlying data generating process.
@@ Acknowledgements
This document provides further information to accompany the main paper. Section S:2 includes derivations of key properties, further description of methods and additional experiments using some of the standard methods for threshold selection, as referenced in Section 2 of the main paper. Section S:3 provides more detailed descriptions of the Wadsworth, (2016) and Northrop et al., (2017) methods that are used in the analyses of the main paper. Section S:4 presents further details of the distributions in Cases 1-4 of Section 6 of the main paper, including quantile derivations. Section S:5 covers the reasoning for the omission of the Varty et al., (2021) approach from the simulation study in the main paper, justifies the choice of the default tuning parameters for the EQD method and discusses the choice of calibration data within the definition of $d_{b}(u)$ in Section 4 of the main text. Section S:6 presents additional simulation experiments, a detailed breakdown of the results outlined in Section 6 of the main paper, and an exploration into the effect of taking different candidate threshold grids. Finally, Section S:7 provides a more extensive analysis of the coverage of true quantiles achieved by both Algorithm 1 and 2, as mentioned in Section 5 of the main text.
@@
$F_{X}(x)=\int_{0}^{x}\frac{h(s;\sigma,\xi)\mathbb{P}(\mathcal{B}<s)}{q+\bar{H}(1;\sigma,\xi)}\ \mathrm{d}s$ (S.6)
where $\bar{H}(x;\sigma,\xi)$ and $h(x;\sigma,\xi)$ are the survivor and density functions of a GPD with parameters $(\sigma,\xi)$ and a threshold 00, $\mathcal{B}\sim\text{Beta}(\alpha,\beta)$, so $0\leq\mathcal{B}\leq 1$ and $q=\int_{0}^{1}h(s;0.5,0.1)\mathbb{P}(\mathcal{B}<s)\mathrm{d}s$. This unusual distribution function transitions from a non-GPD distribution to an exact GPD for excesses of $1$ (the upper bound of $\mathcal{B}$), so that the true threshold for $X$ is $u=1$ and the excess distribution, for $x>1$, has survivor function $\bar{H}(x-1;\sigma_{1}=\sigma+\xi,\xi)$ following from the threshold stability property (S.1). Here $\tau:=\mathbb{P}(X\leq 1)=q/(q+\bar{H}(1;\sigma,\xi))$ so the true quantile $x_{p}$ for $X$ with exceedance probability $p$, where $p\leq 1-\tau$, is:
@@
Although $F_{X}$ appears complex, it is straightforward to simulate samples from $X$. Specifically, $X$ is generated using a rejection method by first simulating a proposal variable $Y\sim\text{GPD}(\sigma,\xi)$ above the threshold of 00. The rejection step involves rejecting $Y$ as a candidate for $X$ only if $Y<\mathcal{B}$, where $\mathcal{B}$ is a random variate generated from a $\text{Beta}(\alpha,\beta)$ distribution. Since $\mathcal{B}\leq 1$, for $Y>1$ it follows that $X=Y$, but only a proportion of the values of $Y<1$ are retained in $X$, with the rate of retention dependent on the parameters of the $\text{Beta}(\alpha,\beta)$ distribution.
@@
To further emphasise the differences between Cases 1-4, Figure S.4 shows return level plots for the simulated distributions of Cases 1-4 for a range of return periods (per observation). In particular, this plot emphasises the key difference between Case 3 and the other cases; Case 3 has a finite upper-endpoint due to $\xi<0$ whereas the other cases have unbounded distributions. This difference is not apparent from the density plots in Figure S.3.
@@
Figure S.4: True return values of simulated datasets from Cases 1-4.
@@
In this section, we provide evidence, based on a range of simulation studies, to support several decisions we made in Section 4 of the main paper. Specifically, in Section S:5.2, we provide evidence to demonstrate the advantages of using the EQD over the Varty et al., (2021) method; Section S:5.3 outlines results indicating that our suggested default choices for the tuning parameters $(B,m)$ of the EQD method are widely suitable, and that there is very little sensitivity to these choices; in Section S:5.4, we show that the choice of bootstrap data as the calibration data in the metric $d_{b}(u)$ works comparably relative to using the observed data; and in Section S:5.5, we demonstrate the benefits of the bootstrapping component of the EQD method.
@@
We assess the performance of the two methods based on both the selection of thresholds (as the truth is known in each case) and on the subsequent estimation of quantiles for a range of exceedance probabilities, namely the $(1-p_{j,n})$-quantiles where $p_{j,n}=1/(10^{j}n)$ for $j=0,1,2$ with $n$ denoting the length of the simulated dataset. For each case and each of the measures of fit, all comparisons between the methods are based on estimates obtained using the same set of 500 replicated samples, so any difference in the methods found is simply due to the two method’s performance.
@@
0 0.190 0.190 $\bm{0.677}$ 0.705
1 $\bm{0.323}$ 0.324 $\bm{1.563}$ 1.673
2 0.483 0.483 $\bm{3.043}$ 3.378
Table S.2: RMSE of the estimated $(1-p_{j,n})$-quantiles in Cases 1-4 based on fitted GPD above chosen threshold for the EQD and Varty et al., (2021) methods. The smallest value for each quantile are highlighted in bold.
@@
0 0.036 0.036 $\bm{0.379}$ 0.444
1 0.105 0.105 $\bm{1.926}$ 2.479
2 0.233 0.233 $\bm{7.287}$ 10.299
Table S.3: Bias [left] and variance [right] of the estimated $(1-p_{j,n})$-quantiles in Cases 1-4 based on fitted GPD above chosen threshold for the EQD and Varty et al., (2021) methods. The smallest variance and absolute bias for each quantile are highlighted in bold.
@@
This section provides the sources of evidence presented in the main paper for Sections 4.2 and 4.3 in terms of the effect of $m$ on the interpolation of quantiles in the metric and the suitability of our default values of the tuning parameters $(B,m)$.
@@
First, consider an analysis of the sensitivity of the EQD method to different choices of its two tuning parameters $(B,m)$, where $B$ denotes the number of the bootstraps for which $d_{b}(u)$ is evaluated in order to calculate the overall metric $d_{E}(u)$, and $m$ is the number of quantiles used in the evaluation of the metric $d_{b}(u)$ for each bootstrap. In the main paper, the values of $(B,m)=(100,500)$ are proposed as the default values for the simulation studies of the performance of the EQD method. These values are used for the tuning parameters in all simulation studies of the EQD method in the main paper and the supplementary material.
@@
We focus our sensitivity analysis on the 500 replicated samples of Case 1 where $n=1000$, detailed in Section 6 of the main paper. Tables S.4 & S.5 provide the RMSEs of the threshold estimates along with the computation time relative to that of the default value when using the EQD with different values of $B$ and $m$ respectively.
@@
For the choice of $B$, the number of the bootstraps for which $d_{b}(u)$ is evaluated when calculating $d_{E}(u)$, Table S.4 shows that the computation time of the EQD method increases linearly with $B$. As $B$ is simply the number of bootstrap samples in an average, then in principle we want to take $B$ as large as possible to remove Monte Carlo noise in the average approximation of the associated expectation. Thus, in selecting $B$, we require it to be sufficiently large so that any residual Monte Carlo noise is not important (for threshold selection to be stable) whilst recognising the linear increase in computation time from this choice. Thus, for one-off analyses, as computation time is not of particular concern, it is ideal to take $B$ as large as possible. However, for simulation studies, a more careful choice of $B$ is required as accuracy needs to be balanced with computation time. Table S.4 provides evidence on how the Monte Carlo noise is diminishing as $B$ increases, and shows the RMSE value stabilising as $B$ increases. The decreases in RMSE are only very slight, especially when compared to the differences we see between the EQD and other existing approaches in Section 6 of the main paper.
$B$ 200 400 1000 100 (default)
@@
Table S.4: RMSEs for threshold estimates and for the relative computation time compared to the default choice of $B=100$ obtained using the EQD method for different values of $B$ for Case 1. Each result uses $m=500$ and 500 replicated samples.
For the choice of $m$, the number of quantiles used in the evaluation of the metric $d_{b}(u)$ for each bootstrap, Table S.5 shows results for two different strategies for selecting $m$: the first allowing $m$ to be proportional to the data sample size $n$ irrespective of threshold value, i.e. $m=cn$ for $c=0.5,1,2,10$, and the second allowing $m$ to vary according to the number, $n_{u}$, of exceedances of each candidate threshold $u$, i.e. $m=cn_{u}$ for $c=0.5,1,2,10$. Our reason for exploring the second strategy is to ensure that across candidate thresholds we are using the same level of interpolation/extrapolation to non-sample quantiles. For each strategy we examine the effect of different degrees of proportionality $c$. The RMSE of the threshold estimates obtained using the EQD show little sensitivity to the value of $m$ across the two strategies and all levels of proportionality.
@@
Table S.5: RMSEs for threshold estimates and for the relative computation time compared to the default of $m=500$ of the EQD method for different values of $m$ for Case 1. Each result uses $B=100$ and 500 replicated samples.
Table S.5 also shows that increasing $m$ in either strategy is essentially wasting the increased computation time. When estimating $d_{b}(u)$ for a particular bootstrap, we are aiming to approximate the integrated absolute error (IAE) between the model quantiles and sample quantiles for that sample. This $d_{b}(u)$ then feeds into the overall $d_{E}(u)$ for the excesses of candidate threshold $u$ which approximates the IAE between model quantiles and the data generating process. This sensitivity analysis shows that once we choose a suitably large value for $m$, any changes in this approximation error are quite small in comparison to the $d_{b}(u)$ value itself.
Now consider the effect of $m$ on the interpolation of quantiles in the metric. While we have shown that changing $m$ does not meaningfully affect the RMSE of threshold choice over repeated samples, it is still important to investigate the effect of this choice for the values of $d_{b}(u)$ and $d_{E}(u)$ for a range of candidate thresholds as well as the effect, if any, on the resulting threshold choice for a particular dataset. In particular, we are interested in the effect of the choice of interpolation grid between $m=500$ and $m=n_{u}$.
For a particular bootstrap sample, the choice of $m=500$ can lead to under- or over-sampling of the upper tail when approximating the IAE of the QQ-plot, depending on if $m<n_{u}$ or $m>n_{u}$. While this may not be ideal, it is only important if it has a significant effect on the overall metric value in a way that unfairly or adversely affects the resulting choice of thresholds.
To explore the effect of the interpolation grid on the sampling distribution of metric values for thresholds (i.e., $d_{b}(u)$ values), $d_{E}(u)$ and on the resulting threshold choice, we have considered the following additional investigations. Let $d^{500}_{b}(u)$ and $d^{n_{u}}_{b}(u)$ denote the value of the metric for the $b^{\text{th}}$ bootstrap sample above threshold $u$ using $m=500$ and $m=n_{u}$ respectively. For the first simulated sample of Case 1 and the Gaussian case, we look at:
  1. 1.
    The distribution of $d^{500}_{b}(u)$ and $d^{n_{u}}_{b}(u)$ for each $u$ over the candidate grid of thresholds.
  2. 2.
    The distribution of the relative difference between $d^{500}_{b}(u)$ and $d^{n_{u}}_{b}(u)$ to the overall metric $d_{E}(u)$ over the candidate grid.
The reasons that we selected these two features to investigate are that the former looks at the effect of the interpolation on the sampling distribution of the metric while the latter assesses if the interpolation method could significantly change the overall metric value for any particular thresholds on a particular dataset.
For the first simulated data sample of Case 1, we explore the distribution of $d^{500}_{b}(u)$ and $d^{n_{u}}_{b}(u)$ in Figure S.5. Specifically, this figure shows boxplots of the distribution of $d^{500}_{b}(u)$ and $d^{n_{u}}_{b}(u)$ values for each value of $u$. The mean of these values, i.e., $d_{E}(u)$, is also plotted as a black point in each boxplot. For particular thresholds, comparison of the sampling distributions of $d^{500}_{b}(u)$ and $d^{n_{u}}_{b}(u)$ values shows only very minor differences. While there are some larger differences between the plots, particularly at higher thresholds, the black points in each plot indicate that these differences in $d_{b}(u)$ do not lead to large differences in the overall metric value $d_{E}(u)$.
Figure S.5: Boxplots of $d^{500}_{b}(u)$ [left] and $d^{n_{u}}_{b}(u)$ [right] for each $u$ over the candidate grid of thresholds for the first sample of Case 1. The mean for each threshold is shown as a black point.
To demonstrate our findings from Figure S.5 more concretely, Figure S.6 shows the sampling distribution of the difference $d^{500}_{b}(u)-d^{n_{u}}_{b}(u)$ relative to the metric value $d_{E}(u)$ for each threshold $u$. Across almost all thresholds, most of the values within the bootstrap sampling distribution lie very close to zero and more importantly, the mean of the sampling distribution lies very close to zero. These findings indicate that the choice of interpolation grid has no meaningful effect on the value of the metric for a particular threshold, in general. However, for the very largest threshold shown in the plot, we see a much larger range for the bootstrapped distribution of these relative differences, showing the effect of the over-sampling of the upper tail as you mentioned in your original review. Importantly, referring back to Figure S.5, the $d_{b}(u)$ values for the highest threshold are all larger than the largest $d_{b}(u)$ value for the optimal threshold in both plots. Thus, while there is a clear effect on the values of $d^{500}(u)$ and $d^{n_{u}}(u)$ for particular bootstrap samples, both Figure S.5 and S.6 show that the effect on the mean is relatively small and certainly, would not alter the selected threshold in any way.
Figure S.6: Plot of $(d^{500}_{b}(u)-d_{b}^{n_{u}}(u))/d_{E}(u)$ for first sample of Case 1. The mean for each threshold is shown as black points.
For Case 1, the true threshold is at a low sample quantile (16.67%) and so, it is unlikely that the threshold choice would be affected by the under- or over-sampling of the tail as the only significant effect of this comes at very high thresholds. In a case where the optimal threshold lies at a higher sample quantile, based on the above analysis, we might expect the threshold choice to show greater sensitivity to choice of interpolation grid. To explore a case of this nature, we repeat the above analysis on the first sample from the replicated data of the Gaussian case. Here, we expect the optimal threshold to lie further into the tail due to the slow convergence of the Gaussian distribution to an extreme value limit. As a result, in theory, we expect that this case could show more sensitivity to the under- or over-sampling of the upper tail.
Figures S.7 and S.8 show the Gaussian results, in the same format as for Case 1. In Figure S.7, there is a clear threshold choice in both plots and there does not seem to be any clear effect from the choice of interpolation grid on the bootstrap sampling distribution of mean-absolute deviations, and certainly there is no clear effect on the mean values for any thresholds. In fact, surprisingly, any effect due to under- or over-sampling the upper tail is much smaller in this case than above. Figure S.8 reiterates this where the sampling distribution of relative differences lie close to zero for all thresholds. The choice of interpolation grid does not show any meaningful effect on the overall metric value and certainly, would not affect the choice of threshold for this dataset.
Figure S.7: Boxplots of $d_{b}(u)$ for the first sample of the Gaussian case with $m=500$ (left) and $m=n_{u}$ (right). The mean for each threshold is shown as black points.
Figure S.8: Plot of $(d^{500}_{b}(u)-d_{b}^{n_{u}}(u))/d_{E}(u)$ for first sample of the Gaussian case. The mean for each threshold is shown as black points.
Given that our overall aim is to approximate the IAE between model quantiles and quantiles of the data generating process, we choose to keep the quality of the approximation of the IAE for particular samples fixed for a fair comparison across a range of candidate thresholds. Since $m$ controls the quality of the approximation, we suggest $m=500$ is sufficient based on the above results. Given we are confident in the quality of results for this choice, for convenience, we keep this fixed value across sample sizes and threshold levels in our simulation studies.
In conclusion, in order to balance accuracy and computation time for our large scale simulation studies, we find that the EQD method, with value $B=100$ and $m=500$, provides sufficiently accurate results in a timely manner, so we choose to use these value throughout the paper, unless stated otherwise.
@@
Here, we provide results for an adjusted version of the EQD method (as suggested by a referee) with the findings being reported in the main paper in Section 4.2. Here, the adjusted version takes the calibration data used for threshold estimation as the actual observed sample excesses $\bm{x}_{u}$ of candidate threshold $u$, as described below. This adjusted version differs from the EQD method as proposed in Section 4.1 of the main paper, where the calibration data for a particular bootstrapped sample are calculated based on the bootstrapped excesses $\bm{x}^{b}_{u}$. Thus, the only difference between the two approaches is that in $d_{b}(u)$, given by metric (2), the $Q(p_{j};\bm{x}^{b}_{u},\bm{q})$ term is replaced by $Q(p_{j};\bm{x}_{u},\bm{q})$ in the adjusted version.